OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

do you know how to do point slope?

OpenStudy (anonymous):

A little. If I am given an equation I can solve it but I am not sure what they are asking in this.

OpenStudy (zale101):

use the equation y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

y - 10 = 3x (x - 1)

OpenStudy (anonymous):

if you use the quation m= y2-y1 over x2-x1 you will get the slope

OpenStudy (zale101):

does this formula seem familiar to u y=mx+b?

OpenStudy (zale101):

@MiaS

OpenStudy (anonymous):

I know y = mx + b but not @melody16 s

OpenStudy (zale101):

Can you plug in the numbers for me? (-2, 1) and (1, 10)

OpenStudy (zale101):

Do you know how to do that?

OpenStudy (anonymous):

To find the equation we have : (-2, 1) and (1, 10)? \[y=mx+b\] where m is the slope. We can calculate m by : \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{10-1}{1-(-2)}=3\] To find : b since (1,10) is from the line then : \[10=3\times 1+b\Rightarrow b=10-3=7\] So : \[y=3x+7\]

OpenStudy (anonymous):

What numbers do I plug in? Do I find slope on graph paper first to get y intercept?

OpenStudy (anonymous):

or you could pick a coordinate and plug in the x and y in the slope intercept form to get b

OpenStudy (zale101):

@Noura11 did everything

OpenStudy (anonymous):

Is my answer A

OpenStudy (anonymous):

and what does V ∩ (~H ∩ ~ M) mean. I do not recognize it but it is a question

OpenStudy (anonymous):

What is your last equation exactly ?

OpenStudy (anonymous):

Also. is y = mx + b slope intercept form

OpenStudy (anonymous):

Campbell Music Academy held workshops on classical music composers Mozart, Vivaldi, and Haydn. The Venn diagram shows the number of seniors who chose to attend workshops on each of the three composers. If the total number of seniors in the academy is 64, what is the total number of students who belong to the set V ∩ (~H ∩ ~ M)? 45 33 19 15

OpenStudy (anonymous):

well, b is the y-intercept m is the slope y = mx + b has both of these in it hence, it is slope-intercept form ;)

OpenStudy (anonymous):

Thanks, @funinabox

OpenStudy (anonymous):

y=mx+b slope intercept ax+bx = c standard form y-y1 = m (x-x1) point slope form

OpenStudy (anonymous):



OpenStudy (anonymous):

So what Is it that they are asking. I do not understand it with three numbers. Do I need whatever is not in h or m and only v

OpenStudy (anonymous):

You want anything that is NOT in H, NOT in M, but IS in V.

OpenStudy (anonymous):

so 15

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

Thanks